# Automatically Interpreting Millions of Features in Large Language Models

### v1

@@ 3.5.1 Detection
One simple approach for scoring explanations is to ask a language model to identify whether a whole sequence activates a SAE latent given an explanation. Currently, the bottleneck of scoring is on the autoregressive nature of token generation. Detection requires few output tokens for each example used in scoring, meaning examples from a wider distribution of latent activation strengths can be used at the same expense. By including non-activating contexts, this method measures both the precision and recall of the explanation.
Detection is more “forgiving” than simulation insofar as the scorer does not need to localize the feature to a particular token.
@@ 3.5.2 Fuzzing
Fuzzing is similar to detection, but at the level of individual tokens. Here, potentially activating tokens are delimited in each example and the language model is prompted to identify which of the sentences are correctly marked. Evaluating an explanation on both detection and fuzzing can identify whether a model is classifying examples for the correct reason.
While in detection, non-activating examples were used, in fuzzing we choose activating contexts and randomly select non-activating tokens. Details on the prompt on Appendix A.3.2
@@ 3.5.3 Surprisal
Surprisal scoring is based on the idea that a good explanation should help a base language model $\mathcal{M}$ (the “scorer”) achieve lower cross-entropy loss on activating contexts than it would without a relevant explanation. Specifically, for each context $\mathbf{x}$, activating or non-activating, we measure the information value of an explanation $\mathbf{z}$ as $\log p_{\mathcal{M}}(\mathbf{x}|\mathbf{z})-\log p_{\mathcal{M}}(\mathbf{x}|\tilde{\mathbf{z}})$, where $\tilde{\mathbf{z}}$ is a fixed pseudo-explanation. A good explanation should have higher information value on activating examples than on non-activating ones. The overall surprisal score of the explanation is given by the AUROC of its information value when viewed as a classifier distinguishing activating from non-activating contexts. Details on the prompt and how to compute the score in Appendix A.3.3.
@@ 3.5.4 Embedding
We can imagine explanations of latents as “queries” that should be able to retrieve contexts where the latent is active. These contexts can act as “documents” that are embedded by an encoding transformer, and we can use the similarities between the query and the documents to distinguish between activating and non-activating contexts. If the encoding model is small enough, this technique is the fastest and opens up the possibility to evaluate a larger fraction of the activation distribution. We have seen that using a larger embedding model didn’t significantly improve the scores, see fig A2, although we believe that this approach was under-investigated. Details on the prompt, on the embedding model and on the way to compute the score in Appendix A.3.4.
@@ 3.5.5 Intervention scoring
Unlike the above four context-based scores, intervention scoring interprets a feature’s counterfactual impact on model output. We quantify the interpretability of an intervention $I$ with explanation $\mathbf{z}$ on a distribution of prompts $\pi$ as the average decrease in the scorer’s surprisal about the explanation when conditioned on text generated with the intervention.
@@
See Appendix A.7 for details on the explanation and scoring pipeline we use in our intervention experiments.
@@ 4.1 Comparing scoring methods
Since simulation scoring is an established method, we measure how our other context-based scoring techniques correlate with simulation, as well as how they correlate between themselves (see tables 1 and A1).
We find that fuzzing and detection have the highest Spearman correlations with simulation scoring, at 0.70 and 0.44 respectively. The imperfect correlations hint at either shortcomings of the scoring metrics or the fact that these metrics can measure different qualities of explanations.
Simulation correlates more strongly with fuzzing than with detection, which is expected because fuzzing is similar to simulation scoring. It correlates weakly with surprisal and embedding scoring. On the other hand, surprisal and embedding scores correlate more with detection scoring than with fuzzing, and most strongly with each other. Due to the speed of embedding scoring, we believe it to be a potentially scalable scoring technique for quick iteration.
@@ 4.1.1 Intervention scoring
In figure 4, we see that the intervention score we propose is a valuable contribution to the set of automatic interpretability metrics because it (a) distinguishes between features from a trained SAE and random features and (b) recalls features that context-based scoring methods fail to interpret.
Table 1: Spearman correlation computed over 600 different latent scores
Fuzzing Detection Simulation Embedding Surprisal
Fuzzing 1 0.73 0.70 0.41 0.30
Detection 1 0.44 0.71 0.62
Simulation 1 0.33 0.20
Embedding 1 0.79
Surprisal 1
@@
Figure 3: Fuzzing and detection scores for different sampling techniques. Panels a) and b) show the distributions of fuzzing and detection scores, respectively, as a function of different example sampling methods for explanation generation. Sampling only from the top activation gets on average lower accuracy in fuzzing and on detection when compared with randomly sampling and sampling from quantiles. The distributions from random sampling and sampling from quantiles are very similar. Panels c) and d) measure how the explanations generalize across activation quantiles, showing that explanations generated from the top quantiles are better at distinguishing non-activating examples in detection, but have lower accuracy on other quantiles, especially on the lower activating deciles. This also happens with explanations generated from examples sampled randomly and from quantiles, but the accuracy does not drop as much in lower activating deciles.
4.2 Comparing explanation methods
@@
We use a set of 500+ latents as a testbed to measure the effects of design choices and hyperparameters on explanation quality. Each latent is scored using 100 activating and 100 non-activating examples. The activating examples are chosen via stratified sampling such that there are always 10 examples from each of the 10 deciles of the activation distribution.
We evaluate explanation quality using fuzzing, detection and embedding scores, as those were both quick to compute and easy to interpret. We expect this mix of scores to correlate well with simulation scoring and reflect the extent to which the proposed explanation is valid over both activating and non-activating examples.
@@
Table A6 shows that explanations based on more examples tend to have higher scores. By sampling the examples that are shown to the explainer model instead of showing just the top activating examples it is possible to increase the scores of the features, see fig. 3 and A5. This effect would not be seen if the scoring were done on just the most activating examples, underscoring a problem with current auto-interpretability evaluations, which produce explanations using top activating examples and evaluate them on a small subset of the activation distribution.
@@
Increasing the size of the explainer model increases the scores of the explanations but we don’t find the performance of Claude Sonnet 3.5 to be much higher than that of Llama 3.1 70b, see A7, and both are similar to those generated by a human. We expect that this may be due to the fact that the prompting techniques were first optimized for Llama 3.1, and that there could be potential gains in optimizing the prompting technique of Claude. Not surprisingly, we also see that even for fuzzing and detection, using a smaller scorer model leads to lower scores.
In table A4, we don’t find a significant dependence on whether we use the same dataset as the training set of the SAEs (see A2) whether we used chain of thought or not (see A3) and whether we change the size of the contexts shown.
4.3 Comparing SAEs
SAEs with more latents have higher scores, and scores that are significantly higher than those of neurons, which are just slightly better than randomly initialized SAEs, see A8. Neurons are more interpretable if made sparser by only considering the top $k$ most activated neurons on a given token, but still significantly underperform SAEs in our tests, see A8.
@@ 4.3 Comparing SAEs
The location of the SAEs matters; residual stream SAEs have slightly better scores than ones trained on MLP outputs, see A8. We also observe that earlier layers have lower overall scores, but that the scores across model depth remain constant after those layers, see A4.
A detailed analysis of various factors in the auto-interpretability pipeline is in Appendix A.5.
@@
Figure 4: Intervention scores. Here we present intervention scores (Sec 3.5.5) for SAE features in Gemma 2 9B at layer 32. Left: SAE features are more interpretable than random features, especially when intervening more strongly. Our explainer also produces explanations that are scored higher than random explanations. Right: Many features that would normally be uninterpreted when using context-based automatic interpretability are interpretable in terms of their effects on output.
5 Overlap between latents at adjacent layers
Because each block in a transformer performs an incremental additive update to the residual stream, one would expect SAEs trained at nearby layers to learn similar features. Measuring the degree of feature overlap is interesting for a few reasons. First, if adjacent SAEs learn almost identical features, it may not be worthwhile to train and interpret SAEs at every single residual stream layer. Second, we can use feature overlap to sanity check our auto interpretability pipeline: if the explanation for latent $\alpha$ at layer $j$ is very different from the explanation for “the same” latent at layer $j+1$, this would suggest that our pipeline is inconsistent and noisy. Finally, feature overlap may allow us to estimate the degree of semantic similarity between layers, as opposed to mere statistical similarity.
Unfortunately, we cannot simply fix some index $i$ and compare latent $i$ at layer $j$ to latent $i$ at layer $j+1$. This is due to permutation symmetry: given an SAE with parameters $(\mathbf{W}_{e},\mathbf{b}_{e},\mathbf{W}_{d},\mathbf{b}_{d})$ and a permutation matrix $\mathbf{P}\neq\mathbf{I}$, the SAE $(\mathbf{P}\mathbf{W}_{e},\mathbf{P}\mathbf{b}_{e},\mathbf{P}^{T}\mathbf{W}_{d},\mathbf{P}^{T}\mathbf{b}_{d})$ has identical input-output behavior, and yet its latents are completely “unaligned” with the original. Instead, we use the Hungarian algorithm (Crouse, 2016) to compute the permutation which maximizes the Frobenius inner product between the decoder weight matrices of the two SAEs. A similar method was used by Ainsworth et al. (2023) to align weights of independently trained MLPs.
In this section, we focus on the 16k latent SAEs trained on Gemma 2 9b. In Appendix A.6, we report the pairwise Frobenius inner products between SAE decoder matrices at different layers before and after aligning them with the Hungarian algorithm. As expected, this simple metric shows that SAEs trained on adjacent layers of the residual stream are more similar than those trained on the MLPs.
To measure the semantic similarity between the explanations, we embed the explanation of each latent and organize the embeddings into a matrix, sorted with the indices found by the Hungarian algorithm. We then compute the Frobenius inner product of these matrices, ignoring the latents that don’t have an explanation on both layers. We observe that the residual stream SAEs have higher semantic overlap for neighboring layers than the MLP SAEs.
Our results suggest that, when training SAEs given a limited compute budget, one should prioritize training wider SAEs on a smaller subset of residual stream layers, perhaps every $k$th layer for some stride $k>1$. If feature diversity is a priority, it may make more sense to train SAEs on MLP outputs exclusively, although we find MLPs to be slightly less interpretable than the residual stream, see A8.
It’s worth noting that the Hungarian algorithm could also be used to align latents from independently trained SAEs in other contexts, such as SAEs trained on different checkpoints of the same training run, or even completely different architectures trained on the same data (Moschella et al., 2022). We leave these extensions to future work. [^3] All that is needed is an adequate cost function for measuring the similarity of individual latents. SciPy’s implementation supports rectangular cost matrices, allowing SAEs of different sizes to be (partially) aligned.
Figure 5: Alignment between embeddings of explanations of the 16k Gemma 2 9b SAEs. After “aligning” the features of each layers by using the decoder weights, we observe that the explanations found for the latents of the residual stream SAE are more similar than for latents of the MLPs. In this figure, alignment is the measured by the Frobenius inner product between the feature embeddings.
@@ 6 Conclusion
Explaining the latents of SAEs trained on cutting-edge LLMs is a computationally demanding task, requiring scalable methods to both generate explanations and assess their quality. We addressed issues with the conventional simulation-based scoring and introduced four new scoring techniques, each with distinct strengths and limitations. These methods allowed us to explore the “prompt design space” for generating effective explanations and propose practical guidelines.
While most latents are well-explained by contexts of length 32, we expect features active over longer contexts to be important and plan to develop better methods for capturing these long-context features. Additionally, although current scoring methods do not account for explanation length, we believe shorter explanations are generally more useful and will incorporate this in future metrics. Some scoring methods also require further refinement, particularly in selecting non-activating examples to improve evaluation.
By analyzing a large number of explanations, we examined overlaps across layers and the distribution of explanations in decoder direction space. Our results suggest that, when compute resources are constrained, it may be more efficient to train wider SAEs on a small subset of residual stream layers, rather than narrower SAEs on all layers.
Access to better, automatically generated explanations could play a crucial role in areas like model steering, concept localization, and editing. We hope that our efficient scoring techniques will enable feedback loops to further enhance the quality of explanations.
@@ 7 Data and code availability
In collaboration with Neuronpedia we also release the explanations for the Gemma 2 9b 16k residual stream SAEs, the 16k MLP SAEs, the 131k residual stream SAEs and the Llama 3.1 8b 262k residual stream and MLP SAEs. Explanations can be directly downloaded here.

### v2

@@ 3.3.1 Detection
One simple approach for scoring interpretations is to ask a language model to identify whether a whole sequence activates a SAE latent given an interpretation. Detection requires few output tokens for each example used in scoring, meaning examples from a wider distribution of latent activation strengths can be used at approximately the same expense. By including non-activating contexts, this method measures both the precision and recall of the interpretation.
Detection is more “forgiving” than simulation insofar as the scorer does not need to localize the latent to a particular token. This means that detection evaluates whether an interpretation correctly identifies the types of contexts that a latent is active on, even if the interpretation does not correctly identify the actual token.
@@ 3.3.2 Fuzzing
Fuzzing is similar to detection, but at the level of individual tokens. Here, potentially activating tokens are delimited in each example and the language model is prompted to identify which of the sentences are correctly marked. Fuzzing is the most similar with simulation. Because SAE activations are very sparse, simulation scoring mostly boils down to correctly identifying which of the tokens have non-zero activations. For this reason, fuzzing is the score that most correlates with simulation.
When interpretations focus on which tokens the latent activates, but not on which contexts, they can still score high on fuzzing, but comparatively lower on detection, specially if they are common tokens. Thus, evaluating an interpretation on both detection and fuzzing can identify whether a model is classifying examples for the correct reason. Details on the prompt on Appendix A.4.2.
@@ 3.3.3 Surprisal
Surprisal scoring is based on the idea that a good interpretation should help a base language model $\mathcal{M}$ (the “scorer”) achieve lower cross-entropy loss on activating contexts than it would without a relevant interpretation. Specifically, for each context $\mathbf{x}$, activating or non-activating, we measure the information value of an interpretation $\mathbf{z}$ as $\log p_{\mathcal{M}}(\mathbf{x}|\mathbf{z})-\log p_{\mathcal{M}}(\mathbf{x}|\tilde{\mathbf{z}})$, where $\tilde{\mathbf{z}}$ is a fixed pseudo-interpretation. A good interpretation should have higher information value on activating examples than on non-activating ones.
The overall surprisal score of the interpretation is given by the AUROC of its information value when viewed as a classifier distinguishing activating from non-activating contexts. Details on the prompt and how to compute the score in Appendix A.4.3.
Surprisal relies on the ability of the scorer model to tune its predictions based on the explanation, measuring both the relevance of the context given in the interpretation and the token specificity, but we believe that the current setup could be improved, as this is the score with the least correlation with the others.
@@ 3.3.4 Embedding
Classifying between active and non-active contexts given a certain interpretation can also be seen as using interpretations of latents as “queries” that should be able to retrieve relevant “documents”, contexts where the latent is active, between non-relevant “documents”, non-activating contexts. This way, we take a selection of activating and non-activating contexts that embedded by an encoding transformer, and the similarity between the query and the documents is used as a classifier to distinguish between activating and non-activating contexts, and the score is given by the AUROC.
If the encoding model is small enough - we used a 400M parameter model - this technique is the fastest and opens up the possibility to evaluate a larger fraction of the activation distribution. We have seen that using a larger embedding model - 7B parameter model - didn’t significantly improve the scores, see fig A3, although we believe that this approach was under-investigated. Details on the prompt, on the embedding model and on the way to compute the score in Appendix A.4.4.
@@ 3.3.5 Intervention scoring
Unlike the above four context-based scores, intervention scoring interprets a latent’s counterfactual impact on model output. We quantify the interpretability of an intervention $I$ with interpretation $\mathbf{z}$ on a distribution of prompts $\pi$ as the average decrease in the scorer’s surprisal about the interpretation when conditioned on text generated with the intervention.
@@
See Appendix A.7 for details on the interpretation and scoring pipeline we use in our intervention experiments.
@@ 4.1 Comparing scoring methods
When wanting to evaluate the explanation over a larger number of examples, simpler methods than simulation scoring have to be used (Templeton et al., 2024). Simulation scoring is traditionally done only in activating examples, ignoring whether the proposed explanation correctly handles non-activating examples. Simulation scoring is normally done in a small amount of examples per latent due to its high cost, and this is a significant disadvantage that prompted us to investigate more efficient scoring methods.
In Table 1 we compare the amount of tokens used to score a single latent using 100 different contexts, for both fuzzing, detection and two different simulation methods. We can see that fuzzing and detection are at least 5x cheaper than simulation when using the all-at-once (AAO) method described in Bills et al. (2023), although this method requires access to the log-probabilites of the prompt tokens, which are not accessible from providers of state-of-the-art closed-source models. A researcher wanting to generate simulation scores on the fly, without access to a local model, would have to use the token-by-token (TBT) method, which is 30x more than expensive than both fuzzing and detection. For the same set of intepretations, fuzzing and detection scores given by Llama 70b and Claude Sonnet 3.5 have similar distributions. On the other hand, Claude Sonnet 3.5 produces to higher simulation scores on average than Llama 70b, see Table A9 meaning that we should probably compare the price of fuzzing and detection on Llama 70b to the price of simulation using Claude Sonnet 3.5.
Embedding scoring is even cheaper, requiring 4000 input tokens for 100 different contexts, which at an average of \$0.13 per million tokens would only cost \$50/100k latents.
Table 1: Estimated cost of the different scoring methods. Brackets represent tokens that can be cached, potentially saving costs, although we do not do these calculations here. The number of input and output tokens is computed considering 100 examples show, and using fewer examples would cost proportionally less.
Input tokens Output tokens Cost (\$/100k latents) Llama 70b Cost (\$/100k latents) Claude Sonnet 3.5
Fuzzing 19.6k (14.2k) 249 676 6.2k
Detection 17.0k (11.9k) 240 588 5.5k
Simulation (AAO) 104.9k (87.5k) 5 3.6k 31.5k
Simulation (TBT) 496.9k (451k) 46.7k 18.7k 219.1k
We propose that latent interpretations should be evaluated using more than a single technique when possible, because as discussed in Section 3.3, each of the proposed methods have different failure modes, and a more rich scoring framework can address some of these flaws. Specifically, we find that when computing the correlation of scores computed on the same 800 explanations, we find that while there is a clear positive correlation, some scores disagree – see Appendix A.4.6 and Tables A1 and A2.
We find fuzzing to be the most correlated with simulation, as fuzzing is mostly measuring how well the explanation can help the model predict which tokens have non-zero activation. Due to the sparsity of SAE activations, simulation scoring is doing a very similar test. On the contrary, we find that simulation has a significantly lower correlation with detection, embedding and surprisal scoring, as they more accurately measure whether a context activates a given latent, instead of which token is active in a given context. When comparing the simulation scores provided by Claude 3.5 Sonnet, all correlations have a slight increase, see Appendix A.4.6.
We also measure the correlation between our scoring methods and scores given by humans. Humans were tasked to score how an explanation related to a given context, similar to that done in (Templeton et al., 2024). The human evaluators could choose between 4 options, “1 - Explanation not related to the text”, “2 - Explanation vaguely related to the text”, “3 - Incomplete explanation but present in the text”, “4 - Explanation describes the activation of the tokens”. We measure the average score of a latent in at least 5 contexts. In total 700 contexts and 81 latent interpretation. We find that fuzzing as the best Spearman correlation with human scores (0.69), followed by simulation (0.60), and detection (0.59). Surprisal and embedding have 0.34 and 0.32 Spearman correlation respectively.
Due to their relative cheapness compared with simulation scoring, we propose that fuzzing and detection scoring can be used to estimate whether explanations correctly identify on which tokens a latent is most likely to be active, and in which types of contexts this happens. Cheaper methods like embedding can be used to quickly iterate on explanations or to broadly separate latents that have bad explanations from those with good explanations, so that they can be refined.
@@ 4.1.1 Intervention scoring
The previously discussed scoring methods are correlational in the sense that they measure whether an interpretation can be used to predict the activation values of a given latent, or distinguish between activating and non-activating examples. Intervention scoring proposes to measure how well a given interpretation can predict the effect of interventions on the corresponding latent. Here we compare correlational interpretations generated with our pipeline and scored with fuzzing, to a set of interventional interpretations scored with intervention scoring. Our hypothesis is that some latents will have low correlational scores because their behavior is better explained by their downstream effects than by the contexts where they are active.
In Figure 4, we see that there is a slight negative correlation between the fuzzing score and the intervention score, showing that on average, latents with low fuzzing score can be better explained using their downstream effects than latents with high fuzzing score. We also find that the distribution of intervention scores of interpretation on a trained SAE is significantly different from the distribution of an interpretation of a randomly initialized SAE or a real SAE with randomly assigned interpretations, supporting the validity of this scoring method.
@@
Figure 3: Fuzzing and detection scores for different sampling techniques. Panels a) and b) show the distributions of fuzzing and detection scores, respectively, as a function of different example sampling methods for interpretation generation. Sampling only from the top activation gets on average lower accuracy in fuzzing and on detection when compared with uniform sampling and stratified sampling. The distributions from random sampling and sampling from quantiles are very similar. Panels c) and d) measure how the interpretations generalize across activation quantiles, showing that interpretations generated from the top quantiles are better at distinguishing non-activating examples, but have lower accuracy on other quantiles, especially on the lower activating deciles. This also happens for the other interpretations, but the accuracy does not drop as much in lower activating deciles. We also show the scores of random explanations.
4.2 Comparing interpretation methods
@@
We use a set of 500+ latents as a testbed to measure the effects of design choices and hyperparameters on interpretation quality. Each latent is scored using 100 activating and 100 non-activating examples. The activating examples are chosen via stratified sampling such that there are always 10 examples from each of the 10 deciles of the activation distribution. We evaluate interpretation quality using fuzzing, detection and embedding scores, as those were both quick to compute and easy to interpret. We expect this mix of scores reflect the extent to which the proposed interpretation is valid over both activating and non-activating examples.
@@
We find that the interpretations generated by considering only the activating contexts are significantly different from those generated by selecting activating contexts from the whole distribution. Interpretations based on top examples have higher specificity, but lower sensitivity. In fact, we observe that the sensitivity depends on how the degree of activation of each example shown, and that this sensitivity decreases faster for interpretation based on top examples. Interpretations generated from top activating examples are more concise and specific interpretations but fail to capture the whole distribution, while explanations based on examples sampled from the whole distribution lead to interpretations that are too broad, see Appendix A.3. This effect would not be seen if the scoring were done on just the most activating examples, underscoring a problem with current auto-interpretability evaluations, which produce interpretations using top activating examples and evaluate them on a small subset of the activation distribution.
@@
Increasing the size of the explainer model increases the scores of the interpretations, but we don’t find the explanations generated by Claude Sonnet 3.5 to have much higher scores than those generated by Llama 3.1 70b, see A9, and both are similar to those generated by a human. Not surprisingly, we also see that even for simpler scores like fuzzing and detection, using a smaller scorer model leads to lower scores. It is possible that the explanations generated by Claude would be better than those generated by Llama 3.1 70b had we optimized the prompting techniques for that model.
Showing the explainer model a larger number of examples leads to slightly higher scores (Table A7). We find it doesn’t matter much whether we use the same dataset as the training set of the SAEs (Table A3) or whether we slightly change the size of the contexts shown (Table A5). Using COT, at least when generating interpretations using Llama 70b, does not significantly increase their quality (Table A4) while significantly increasing the compute and time required to generate them. For this reason, we have not used it for our main experiments.
@@
We find that SAEs with more latents have higher scores, and scores that are significantly higher than those of neurons. Neurons are more interpretable if made sparser by only considering the top $k$ most activated neurons on a given token, but still significantly underperform SAEs in our tests (Table A10). The location of the SAE matters; residual stream SAEs have slightly better scores than ones trained on MLP outputs (Table A10). We also observe that earlier layers have lower overall scores, but that the scores across model depth remain constant after the first few layers (Fig A5).
@@
Figure 4: Intervention scores. Here we present intervention scores (Sec 3.3.5) for SAE latents in Gemma 2 9B at layer 32. Left: SAE latents are more interpretable than random latents, especially when intervening more strongly. Our explainer also produces interpretations that are scored higher than random interpretations. Right: Many latents that would normally be uninterpreted when using context-based automatic interpretability are interpretable in terms of their effects on output.
@@ 5 Conclusion
Explaining the latents of SAEs trained on cutting-edge LLMs is a computationally demanding task, requiring scalable methods to both generate interpretations and assess their quality. We addressed issues with the conventional simulation-based scoring and introduced five new scoring techniques, each with distinct strengths and limitations. These methods allowed us to explore the “prompt design space” for generating effective interpretations and propose practical guidelines.
Additionally, although current scoring methods do not account for interpretation length, we believe shorter interpretations are generally more useful and will incorporate this in future metrics. Some scoring methods also require further refinement, particularly in selecting non-activating examples to improve evaluation. We also investigated a method to generate and scoring interpretations of latents that is based on their downstream effect and shown that low score “correlational” scores could be due to the existence of “output” features.
Access to better, automatically generated interpretations could play a crucial role in areas like model steering, concept localization, and editing. We hope that our efficient scoring techniques will enable feedback loops to further enhance the quality of interpretations.
@@ 6 Data and code availability
In collaboration with Neuronpedia we also release the interpretations for the Gemma 2 9b 16k residual stream SAEs, the 16k MLP SAEs, the 131k residual stream SAEs and the Llama 3.1 8b 262k residual stream and MLP SAEs. interpretations can be directly downloaded here.